A firm has the following financial statement info:

Sales $        6,000,000
COGS $        5,400,000
Cash $              50,000
Inv $            550,000
AR $            390,000
AP $            450,000
Accruals $              60,000
NP $            400,000

For each question, round to 2 decimals.

What is the length of its inventory conversion period?   

What is the length of its average collection period?

What is the length of its payables deferral period?

What is the length of its cash conversion cycle?
Answer #1

1

Inventory turnover = COGS/inventory
Inventory turnover = 5400000/550000
Inventory turnover = 9.82
days of inventory on hand = number of days in a year/inventory turnover
days of inventory on hand = 365/9.82
days of inventory on hand = 37.17
2
Receivables turnover = Credit sales/receivables
Receivables turnover = 6000000/390000
Receivables turnover = 15.38
days of sales outstanding = number of days in a year/receivables turnover
days of sales outstanding = 365/15.38
days of sales outstanding = 23.73
3
Accounts payables turnover = purchases/payables
Accounts payables turnover = 5400000/450000
Accounts payables turnover = 12
days of payables outstanding = number of days in a year/accounts payable turnover
days of payables outstanding = 365/12
days of payables outstanding = 30.42
4
Operating cycle = days of sales outstanding + days of inventory on hand
Operating cycle = 23.73+37.17
Operating cycle = 60.9
Cash conversion cycle = Operating cycle - days of payables outstanding
Cash cycle = 60.9-30.42
Cash cycle = 30.48